# Constant search time algorithm via topological quantum walks

### v1

@@ Abstract
The change of time asymptotic that is shown in Fig.2(b) happens when the system size becomes larger that diameters of both trapped state pairs. In Supplemental material [21] we show that the state from the pair with smaller overlap in Fig.2(c) has larger radius oflocalization ($95\%$ of the probability density). This pair of states determines the chance in the search time asymptotic as shown in Fig. 2b.
Finally, we note on stability of algorithm in comparison with Grover’s search. In Supplemental material [21] we present a comparison from Fig.2(d) for the disordered systems with disorder being implemented in rotation angles. In the large disorder cases walker ‘2’ shows similar or even better performance that Grover’s search implemented via QRW, and is also advantageous against classical search.
In conclusion, we demonstrated that a fast quantum search algorithm can be realized in topological split-step quantum random walks. We proposed a concrete implementation of this search consisting of the following steps, namely (i) Check the appearance of pair of trapped states near a defect and overlap product maximization criteria for a small system; (ii) apply the same encoding of searched site in the large system with the same walker parameters;(iii) initialize uniform state in the large system; (iv) the high density peak in time will reveal the searched node position [24]. Our algorithm admittedly relies on the proper tuning of walker and searched node parameters. At the same time this property also allows for a protection against disorder with $\Theta_{1}$ parameter values being not within efficient search interval. Interestingly, in the search algorithm we present here the ‘structure’ is induced in a quantum way (via dynamically appearing interference pattern due to overlap with trapped states) after fixed amount of time, i.e. it is not known a-priori where the pairs of trapped states are localized. In such formulation, our algorithm is suitable to make fast detectors of emergent defects in topological split-step quantum random walks. In addition, we point out that the dependence of algorithm of superposition of only four states makes it less sensitive to disorder rather that Grover’s algorithm, which performs optimally when a significant percentage of states contributed into constructive interference. Our search algorithm, as formulated in this work, can be immediately tested in photonic and synthetic lattice quantum walk experiments that realize topological QRW [25, 26, 27, 28].
@@
Acknowledgements. We benefited from discussions with Anton Akhmerov, Yaroslav Herasymenko, Margarita Davydova and Carlo Beenakker. This project was supported by the Kavli Foundation. This publication is part of the project Engineered Topological Quantum Networks (with Project No. VI.Veni.212.278) of the research program NWO Talent Programme Veni Science domain 2021 which is financed by the Dutch Research Council (NWO). GJ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW).
Supplemental material for “Constant search time algorithm via topological quantum walks”
by D. O. Oriekhov, Guliuxin Jin, and Eliska Greplova
Appendix S1 Properties of split-step quantum random walk
@@
$\lambda^{2}+2\lambda\cos\left(\Theta_{2}/2\right)\left(\sin^{2}\left(\Theta_{1}/2\right)\cos\left(2k_{y}\right)-\cos^{2}\left(\Theta_{1}/2\right)\cos\left(k_{x}\right)\right)+\lambda\sin\left(\Theta_{1}\right)\sin\left(\Theta_{2}/2\right)\left(\cos\left(k_{x}-2k_{y}\right)+1\right)+1=0.$ (S1)
@@
$b=2\cos\left(\Theta_{2}/2\right)\left(\sin^{2}\left(\Theta_{1}/2\right)\cos\left(2k_{y}\right)-\cos^{2}\left(\Theta_{1}/2\right)\cos\left(k_{x}\right)\right)+\sin\left(\Theta_{1}\right)\sin\left(\Theta_{2}/2\right)\left(\cos\left(k_{x}-2k_{y}\right)+1\right)$ (S2)
@@
The topological phases and corresponding Chern numbers were defined in Ref.[8]. The phase separation lines are characterized by the gap closing. For the technical purposed of identifying different types of search speed up discussed in the main text, we plot a map of gap value: the minimal gap value between two bands at $E=0$ or $E=\pm\pi$ (see Fig.S1(a) ) and the maximal gap value around the same energies. From such maps we find that the full gap closing happens only at lines $\Theta_{2}=2\pi n$ with $n$ being integer.
$(a)$Refer to caption $(b)$Refer to caption
@@
Figure S1: The regions of one gap closing in panel (a) and both gaps closing in panel (b). Note the agreement between panel (a) and phase separation lines position.
@@
The colormap of search dependence on defect parameters for the walker ‘1’ shown in Fig.2 of the main text is presented in Fig.S2. The dependence on $\Theta_{2}^{def}$ is absent, which motivates the studying of only different $\Theta_{1}^{def}$ values in the main text.
@@
Figure S2: @D colormap of dependence on defect parameters for a walker ‘1’. The optimal search is performed on a line - which demonstrates that the dependence on $\Theta_{2}^{def}$ is absent.
@@
Appendix S2 The comparison of product of overlaps with optimal search regions
@@
In this section we analyze how precise the correspondence between maximization of $|<d|n><n|i>|^{2}$ and $(\sum_{n=max_{1},max_{2}}|<d|n><n|i>|)^{2}$ is compared to the regions of optimal search in parameter space for a fixed defect parameters. For this purpose we compare the max search probability curves plotted along two lines depicted as dashed cyan in the first panel of Fig.S3. One of these lines is a phase separation line and second one is parallel to phase separation and contains a walker `$1^{\prime}$ discussed in the main text. This second line crosses different topological phases. In the second and third panels of Fig.S3 we plot the comparison of max probability found on time interval $T=1000$ for the system size with $\sqrt{N}=40$ with values of different overlaps. The second and third panels in Fig.S3 clearly demonstrate that the combination of maximum and second maximum products of overlaps - the contribution of two pairs of trapped states - describes the probability peak most precisely. This leads to a conclusion that efficient search is performed by two pairs of trapped states in this system in all cases studied.
@@
Figure S3: Comparison of maximal product of overlaps described in the main text found over all eigenstates of evolution operator $\bar{U}$ for a two lines in parameter space plotted as dashed cyan lines in the first panel. The second panel corresponds to a line going through walker `$1^{\prime}$ and the third panel corresponds to a phase separation line.
@@
Next, we analyze the structure of trapped states and their position in spectrum for the example walker parameters labeled as `$2^{\prime}$. The analysis is similar to what is done in the main text and is presented in three panels of Fig.S4: comparison of search probability with overlap criteria in the first panel, the set of overlaps for each eigenstate of $\bar{U}$ in second panel and the comparison of probability evolution at the defect with the one defined by the two pairs of the trapped states in the third panel. However, in this case we find that the two pairs of trapped states that mainly contribute to the efficient search are placed near zero energy. Also there is no gap in spectrum around this energy, but the in-gap trapped states placed around $E=\pm\pi$ do not contribute into the constructive interference relevant for successful search.
@@
Figure S4: (a) The comparison of search probability with overlap criteria for walker ‘2’. (b) The identification of states that mainly contribute into constructive superposition on a searched node for walker ‘2’ parameters. (c) The comparison of density evolution at searched defect node (blue solid line) with the superposition (red dashed line) of trapped states identified from panel (b).
Appendix S3 Structure of trapped states in coordinate space
@@
In this section we analyze how extended are the trapped states that mainly contribute to the search algorithm. From the search time dependence on system size shown in Fig.2(b) in the main text, we conclude that the more extended trapped state has radius $R\approx 50$ for walker ‘1’ and $R\approx 100$ for walker ‘2’ and defect parameter $\Theta_{1}^{def}=5\pi/8.$. The radius of trapped state was extimated as twice smaller that the system size where transition in time asymptotic happens. To further analyze the trapped state radius dependence on system size, we plot it in Fig.S5. The set of tested system sizes covers the region where the trapped states are having overlap with itself due to periodic boundary condition. To reduce the numerical precision issue related to identifying density of quickly decaying state far from its center, we plot a dependence of radius which covers $80\%$ and $95\%$ of density. Two plots show results for a state taken from each pair of trapped states discussed in the main text in Fig.3(b). The states from second pair have larger radius and it stronger depends on system size. Thus, the change in search time asymptotic is mostly defined by this second pair of trapped states, that has second maximal overlap value.
@@
Figure S5: Radius of trapped state as a function of system size. Left panel corresponds to a trapped state with largest overlap $|\langle d|n\rangle\langle n|i\rangle$ in Fig.3(b) of the main text, and right panel corresponds to the state with second largest overlap.
Appendix S4 Stability of search algorithm with disorder
In this section we make a stability comparison of the proposed algorithm of search based on trapped states with Grover’s algorithm in split-step quantum random walk described in Ref.[7]. The size of the systems is taken to be the same - $512\times 512$ nodes, well above the saturation time threshold for both example walker parameters described in the main text. Since the Grover’s algorithm relies on the coherent superposition of most states present in the system, one could expect that disorder break such constructive superposition faster that for a few trapped states. To check that, we implement the angle disorder of spin rotation matrices on each node. For equation (2) in the main text such disorder is introduced as
$R(\Theta)\to R(\Theta+\delta\Theta_{xy}),\quad\delta\Theta\in[-\delta\Theta_{max},\delta\Theta_{max}],$ (S3)
where the distribution of $\delta\Theta_{xy}$ is uniform in the interval $[-\delta\Theta_{max},\delta\Theta_{max}]$. Similarly, for the split-step QRW analyzed in Ref.[7] and used here as Grover’s algorithm benchmark, the disorder is implemented as
$C_{x}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\\ i&1\end{pmatrix}\to C_{x}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\exp(i\delta\Theta)\\ i\exp(-i\delta\Theta)&1\end{pmatrix},$ (S4)
$C_{y}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\to C_{y}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\exp(i\delta\Theta)\\ -i\exp(-i\delta\Theta)&1\end{pmatrix},$ (S5)
The results of simulation for equal intervals of disorder angle are presented in Fig.S6.
The role of disorder in reducing efficiency of search algorithms in visible in all plots. However, for one of the two walkers, the search probability is always staying higher that the classical probability on the first period of oscillation. Notably, the probability of Grover’s search drops dramatically. In the presented example of disorder configuration and type, the trapped state are demonstrating relatively stable behavior as a basis of search algorithm, even outperforming Grover’s example for large disorder. Still, we point out that the more systematic study should be made to define the best performance conditions of different suboptimal quantum algorithms and compared with Grover’s one in noisy quantum setups. Also one should note that these algorithms still outperform classical search, which behaves as $P_{classical}(t)=t/N$ with $t$ being number of check ups of individual nodes.
Figure S6: The search probability calculated for 2 walker parameters analyzed in the main text (see Fig.2) and compared with Grover’s algorithm in split-step QRW from Ref.[7] and with classical search with linearly increasing probability. The system size is $512\times 512$. Disorder strength varies from small in panel (a) $\delta\Theta_{max}=0.05$, to (b) $\delta\Theta_{max}=0.15$, (c) $\delta\Theta_{max}=0.35$, and large in panel (d) $\delta\Theta_{max}=0.5$. The walker $2$ shows best performance at large disorder out of all compared search procedures.

### v3

@@ III Contribution of trapped states to dynamical localization
The change of time asymptotic that is shown in Fig. 2(b) happens when the system size becomes larger that diameters of both trapped state pairs. In Appendix C we show that the state from the pair with smaller overlap in Fig.2(c) has larger radius of localization ($95\%$ of the probability density). This pair of states determines the chance in the search time asymptotic as shown in Fig. 2b.
Figure 5: The localization probability calculated for 2 walker parameters analyzed in the main text (see Fig.2) and compared with the localization probability for QRW from Ref.[10] and with classical search with linearly increasing probability. The system size is $512\times 512$. Disorder strength varies from small in panel (a) $\Theta_{dis}=0.05$, to (b) $\Theta_{dis}=0.15$, (c) $\Theta_{dis}=0.35$, and large in panel (d) $\Theta_{dis}=0.5$. The walker $2$ shows best performance at large disorder out of all compared search procedures.
IV Dynamical localization on defect in the presence of disorder
Figure 6: The dynamical localization phenomena on a single-node defect averaged over $20$ configurations of disorder. The averaged localization in non-topological split-step QRW from Ref.[10] has a higher density of defect node rather than of a topological QRW due to a larger amount of states contributing to maximum localization. The size of the system is $256\times 256$ nodes.
Now we move on to analyzing the stability of localization on single-node defect with disorder. For the topological QRW the disorder is added to rotation operators on the nodes in Eq.(2), via replacement $R(\Theta)\to R(\Theta+\delta\Theta)$ where $\delta\Theta\in[-\Theta_{dis},\Theta_{dis}]$ is a uniformly distributed random value centered around zero in finite interval. The defect node is not disordered. As an example of another mechanism, the localization of non-topological discrete-time QRW from Ref. [10] is studied for the same system sizes. The corresponding time step operator is defined as:
$U=T_{y}C_{y}T_{x}C_{x},\quad T_{i}|\pm\rangle|i\rangle=|\pm\rangle|i\pm 1\rangle.$ (14)
In this case the disorder is implemented as phases in the coin rotation operators:
$C_{x}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&i\exp(i\delta\Theta)\\ i\exp(-i\delta\Theta)&1\end{pmatrix},$ (15)
$C_{y}(\delta\Theta)=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-i\exp(i\delta\Theta)\\ -i\exp(-i\delta\Theta)&1\end{pmatrix}.$ (16)
While different selections of disorder could be made, the one introduced above is motivated by the possible loss of phase coherence on the nodes of large system (described by slightly different phases acquired by passing each node). The sizes of the test systems are taken to be the same - $512\times 512$ nodes for single configuration of disorder, and $256\times 256$ for averaging over $20$ configurations. Both sizes are well above the saturation time threshold for example walker parameters ‘1’ and ‘2’ described in the previous sections.
The results of simulation for different sizes of interval distributions $\delta\Theta\in[-\Theta_{dis},\Theta_{dis}]$ of angle disorder and for single configuration in coordinate space are presented in Fig. 5. The values $\Theta_{dis}$ are equal for topological and non-topological QRWs. The disorder reduces the height of localization probability peak for both quantum walks, and consequently the corresponding efficiency of search algorithms, as it is shown in panels (a)-(d) for $\Theta_{dis}$ varying from $0.05$ to $0.5$.. However, for one of the two topological walkers, the search probability always remains higher that the classical search probability during the first period of oscillation. Notably, the probability of quantum search (non-topological QRW) drops dramatically as the disorder increases.
While the example shown in Fig. 5 constitutes one specific disorder configuration and type, the trapped states demonstrate relatively stable behavior, even locally outperforming quantum search for large disorder. Still, we point out that the more systematic study should be made to define the optimal performance conditions of different quantum and classical search algorithms. One should note, however, that our algorithm still outperforms classical search in localization probability maximum at fixed time. The classical search always behaves as $P_{classical}(t)=t/N$ with $t$ number of check ups of individual nodes.
Finally, the three panels of Fig.6 demonstrate a study of disorder effect on two mechanisms of localization with average over 20 spatial configurations. We show three different values of disorder interval limit $\Theta_{dis}\in\{0.3,0.5,1.0\}$. The small oscillations of the localization probability are not present in these plots, and the curves are more smooth. The effect of reducing the maximum value of $P_{localization}$ is present, especially for large disorder with $\Theta_{dis}=1$ in panel (c). The localization of non-topological QRW demonstrates a higher probability than both example walkers from topological QRW.
V Conclusions
In conclusion, we demonstrated that a fast dynamical localization phenomenon takes place in the split-step discrete-time topological quantum random walks. It was shown that localization mechanism is based on just a two pairs of localized trapped states associated with defect. Each pair of trapped states is particle-hole symmetry in the quasienergy spectrum. The time of highest localization is defined by the gap that separates two positive-energy trapped states. Due to a finite radius of localization, the position of these states in spectrum is not influenced in large enough systems. We should note that the overlap product criteria served as efficient tool to identify these trapped states, capture the evolution of walker and uncover the principle behind distribution of high-localization peaks in parameter space of walker.
The described phenomena open an avenue towards realization of the search algorithms based on topological QRWs as well as further studies on explainability of localization phenomena in quantum random walks. The search mechanism, as formulated in this work, can be immediately tested in photonic and synthetic lattice quantum walk experiments that realize topological QRW [25, 26, 27, 28].
Acknowledgements.
@@
We benefited from discussions with Tereza Vakhtel, Tomas Osterholt, Anton Akhmerov, Yaroslav Herasymenko, Margarita Davydova and Carlo Beenakker. This project was supported by the Kavli Foundation. This publication is part of the project Engineered Topological Quantum Networks (with Project No. VI.Veni.212.278) of the research program NWO Talent Programme Veni Science domain 2021 which is financed by the Dutch Research Council (NWO). GJ acknowledges the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW).
Code and data
The code to reproduce research in this paper can be found at this URL: https://gitlab.com/QMAI/papers/topologicalsearch. The Zenodo repository containing both code and data can be found at this URL: https://zenodo.org/records/14853734.
Appendix A Properties of split-step quantum random walk
@@
$\lambda^{2}+2\lambda\cos\frac{\Theta_{2}}{2}\left(\sin^{2}\frac{\Theta_{1}}{2}\cos\left(2k_{y}\right)-\cos^{2}\frac{\Theta_{1}}{2}\cos\left(k_{x}\right)\right)$
$+\lambda\sin\left(\Theta_{1}\right)\sin\frac{\Theta_{2}}{2}\left(\cos\left(k_{x}-2k_{y}\right)+1\right)+1=0.$ (17)
@@
$b=2\cos\frac{\Theta_{2}}{2}\left(\sin^{2}\frac{\Theta_{1}}{2}\cos\left(2k_{y}\right)-\cos^{2}\frac{\Theta_{1}}{2}\cos\left(k_{x}\right)\right)$
$+\sin\Theta_{1}\sin\frac{\Theta_{2}}{2}\left(\cos\left(k_{x}-2k_{y}\right)+1\right)$ (18)
@@
The topological phases and corresponding Chern numbers were defined in Ref.[11]. The phase separation lines are characterized by the gap closing. For the technical purposed of identifying different types of search speed up discussed in the main text, we plot a map of gap value: the minimal gap value between two bands at $E=0$ or $E=\pm\pi$ (see Fig.7(a) ) and the maximal gap value around the same energies. From such maps we find that the full gap closing happens only at lines $\Theta_{2}=2\pi n$ with $n$ being integer.
@@
Figure 7: The regions of one gap closing (upper panel) and both gaps closing (lower panel). Note the agreement between the upper panel and phase separation lines position in Fig. 2a.
@@
The colormap of search dependence on defect parameters for the walker ‘1’ shown in Fig.2 of the main text is presented in Fig.8. The dependence on $\Theta_{2}^{def}$ is absent, which motivates the studying of only different $\Theta_{1}^{def}$ values in the main text.
@@
Figure 8: 2D color map of dependence on defect parameters for a walker ‘1’. The optimal search is performed on a line - which demonstrates that the dependence on $\Theta_{2}^{def}$ is absent.
@@
Appendix B The comparison of product of overlaps with optimal localization regions
@@
In this Appendix we analyze how precise the correspondence between maximization of $|<d|n><n|i>|^{2}$ and $(\sum_{n=max_{1},max_{2}}|<d|n><n|i>|)^{2}$ is compared to the regions of optimal localization in parameter space for a fixed defect parameters. For this purpose we compare the max search probability curves plotted along two lines depicted as dashed cyan in the first panel of Fig.9. One of these lines is a phase separation line and second one is parallel to phase separation and contains a walker `$1^{\prime}$ discussed in the main text. This second line crosses different topological phases. In the second and third panels of Fig.9 we plot the comparison of max probability found on time interval $T=1000$ for the system size with $\sqrt{N}=40$ with values of different overlaps. The second and third panels in Fig.9 clearly demonstrate that the combination of maximum and second maximum products of overlaps - the contribution of two pairs of trapped states - describes the probability peak most precisely. This leads to a conclusion that efficient search is performed by two pairs of trapped states in this system in all cases studied.
@@
Figure 9: Comparison of maximized over time localization probability with maximal product of overlaps described in the main text found over all eigenstates of evolution operator $\bar{U}$ for a two lines in parameter space plotted as dashed cyan lines in the first panel. The second panel corresponds to a line going through walker `$1^{\prime}$ and the third panel corresponds to a phase separation line.
@@
Next, we analyze the structure of trapped states and their position in spectrum for the example walker parameters labeled as `$2^{\prime}$. The analysis is similar to what is done in the main text and is presented in three panels of Fig.10: comparison of search probability with overlap criteria in the first panel, the set of overlaps for each eigenstate of $\bar{U}$ in second panel and the comparison of probability evolution at the defect with the one defined by the two pairs of the trapped states in the third panel. However, in this case we find that the two pairs of trapped states that mainly contribute to the efficient search are placed near zero energy. Also there is no gap in spectrum around this energy, but the in-gap trapped states placed around $E=\pm\pi$ do not contribute into the constructive interference relevant for successful search.
@@
Figure 10: (a) The comparison of maximized over time localization probability with overlap criteria for walker ‘2’, with all quantities being normalized by their maximum value on the same grid of $\Theta_{1}^{def}$ parameters. (b) The identification of states that mainly contribute into constructive superposition on a searched defect node for walker ‘2’ parameters. (c) The comparison of density evolution at defect node (blue solid line) with the superposition (red dashed line) of trapped states identified from panel (b).
Appendix C Structure of trapped states in coordinate space
@@
In this section we analyze how extended are the trapped states that mainly contribute to the search algorithm. From the search time dependence on system size shown in Fig.2(b) in the main text, we conclude that the more extended trapped state has radius $R\approx 50$ for walker ‘1’ and $R\approx 100$ for walker ‘2’ and defect parameter $\Theta_{1}^{def}=5\pi/8.$. The radius of trapped state was estimated as twice smaller that the system size where transition in time asymptotic happens. To further analyze the trapped state radius dependence on system size, we plot it in Fig.11. The set of tested system sizes covers the region where the trapped states are having overlap with itself due to periodic boundary condition. To reduce the numerical precision issue related to identifying density of quickly decaying state far from its center, we plot a dependence of radius which covers $80\%$ and $95\%$ of density. Two plots show results for a state taken from each pair of trapped states discussed in the main text in Fig.3(b). The states from second pair have larger radius and it stronger depends on system size. Thus, the change in search time asymptotic is mostly defined by this second pair of trapped states, that has second maximal overlap value.
@@
Figure 11: Radius of trapped state as a function of system size. Left panel corresponds to a trapped state with largest overlap $|\langle d|n\rangle\langle n|i\rangle|$ in Fig.3(b) of the main text, and right panel corresponds to the state with second largest overlap. Two curves correspond to the radius of area in which $80\%$ and $95\%$ of state density concentrated.